OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of -1 and passes through the point (-2, 5) A. y + 5 = -1(x – 2) B. y + 2 = 1(x – 5) C. y + 1 = 5 (x + 2) D. y – 5 = -1(x + 2)

OpenStudy (anonymous):

y=mx+b so you can substitute (-2, 5) and the slope of -1 into it: 5= -1(-2) +b solving for b (the y-intercept) 5=2+b 3=b so the equation is y= -1x + 3 none of these answers are right!!?

OpenStudy (anonymous):

Those are the only answers the teacher gave?